40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 3 15:57:52 2014, Steve, Update, Green Locking, ETMX green power ETMXgreen.jpgXgreenOven36C.jpg
    Reply  Tue Feb 4 01:02:03 2014, Koji, Update, Green Locking, ETMX green power Screenshot-Untitled_Window.png
    Reply  Tue Jan 19 17:27:58 2016, gautam, Update, Green Locking, AUX X power investigations X_END_POWER_BUDGET.JPG
Message ID: 9595     Entry time: Tue Feb 4 01:02:03 2014     In reply to: 9592
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: ETMX green power 

Manasa, Steve: Please revisit the Xend oven temperature again.


I found that the X end SLOW control was left on for ~15days. The output of the filter had grown to ~2e7.

This yielded the laser temperature pulled with the maximum output of the DAC.

This was the cause of the power reduction of the X end SHG; phase matching condition was changes as the wavelength of the IR was changed.

Once the SLOW output was reset, the green REFL was reduced from 4000cnt to 1800cnt.

Attachment 1: Screenshot-Untitled_Window.png  11 kB  | Hide | Hide all
Screenshot-Untitled_Window.png
ELOG V3.1.3-